Could you please help me solve x^2+19x+90=0
Answer by kathydots21(7)


x2+19x+90=0
x2+9x+10x+90=0
x(x2x+32⋅xx)+2⋅5(2⋅5x2⋅5+2⋅32⋅52⋅5)=0
answer:
x∈{−9,(−10)}